Find the maximum magnitude of the linear momentum of a photoelectron emitted when light of wavelength 400 nm falls on a metal having work function 25 eV.

Grade:12

1 Answers

Arun
6 years ago
Energy of photoelectron ⇒ ½ mv^2 = hc/λ – hv base 0 = 4.14 * 10^-15 * 3 *10^8/4 * 10^-7 – 2.5 ev = 0.605 ev. We known KE = P2/2m ⇒ P^2 = 2m * KE. P^2 = 2 * 9.1 * 10^-31 * 0.605 * 1.6 * 10^-19 P = 4.197 * 10^-25 kg – m/s